# Efficiently preparing chiral states via fermionic cooling on bosonic quantum hardware

### v1

@@ Abstract
We propose an efficient protocol for preparing low energy states of arbitrary fermionic Hamiltonians on a noisy bosonic quantum simulator. This procedure involves performing adiabatic cooling by coupling the target system with a simulated bath. The bath is periodically monitored in order to extract entropy from the system. By fermionizing the simulated target system and the bath together, we allow individual fermionic excitations of the system to coherently hop to the bath sites. In this way, we achieve a cooling rate linearly proportional to the density of these excitations, despite the fact that they are non-local in terms of the bosonic degrees of freedom of the hardware. In particular, we show that certain topological phases, such as the chiral (non-Abelian) phase of the Kitaev honeycomb model can be prepared efficiently using our protocol. We find that our protocol performs favorably in the presence of noise, making it suitable for execution on near-term quantum devices.
@@ I Introduction
Simulating quantum many-body systems, both in and out of equilibrium, is one of most promising applications of near-term quantum computers Cirac and Zoller (2012); Georgescu et al. (2014). In such tasks, quantum computers offer a vast advantage over their classical counterparts, by virtue of their ability to store the many-body wave functions and perform unitary gates that implement physical time evolution according to arbitrary Hamiltonians, at a polynomial cost in the system size. Such simulations may have practical utility in many fields, such as materials science, high-energy physics, and quantum chemistry. Preparation of a low energy state of a many-body Hamiltonian is an important subroutine for these simulations. Doing this efficiently for arbitrary Hamiltonians is a highly non-trivial task, and is one of the major challenges in the field.
@@
Figure 1: Scheme for simulated cooling of a fermionic system using bosonic qubits. The system’s fermionic degrees of freedom (black) together with a fermionic “bath” (white) are encoded using bosonic qubits such that the two are charged under the same $\mathbb{Z}_{2}$ gauge field. (a) The system is brought to a flux-free state by measuring the flux operators and anihilating defects in pairs by applying a string of spin operators. (b) The system is evolved unitarily with the time-dependent Hamiltonian in Eq. (2), allowing single fermionic excitations to hop coherently from the system to the bath. (c) At the end of the cooling cycle, the hopping between the system and bath sites is turned off. The location of the fermionic excitations in the bath is measured and the measurement outcomes are used to determine the Hamiltonian for the next cycle.
@@
The preparation scheme consists of two steps. In the first step, the system is prepared in the flux-free sector by measuring the flux operators through all plaquettes, and applying unitary operators that correct the fluxes that are detected Dennis et al. (2002). The second step is to prepare the fermionic ground state, whose Chern number is non-zero. The key idea is to introduce auxiliary “bath” degrees of freedom that simulate an effective fermionic bath, coupled to the system’s emergent fermions. Both species of fermions are coupled to the same $\mathbb{Z}_{2}$ gauge field, allowing single-fermion hopping between them. We provide an explicit encoding of the system and bath fermions into the physical bosonic qubits, generalizing the one-dimensional construction of Ref. Kishony et al. (2023) to two dimensions. The bath fermions are then used to extract energy and entropy from the system, by performing simulated cyclic adiabatic cooling Matthies et al. (2022); Kishony et al. (2023) followed by measurements of the auxiliary qubits. The steps of the protocol are illustrated in Fig. 1.
@@
This paper is organized as follows. In Sec. II we introduce our algorithm for preparing the ground state of the KSL model. In Sec. III we analyze the performance of the protocol both with and without noise via numerical simulations. Finally, in Sec. IV we generalize our protocol for arbitrary interacting fermionic systems.
II Preparing a chiral spin liquid
II.1 Model
@@
Here, $\mathbf{J}=(i,j,s)$ denotes the sites of the honeycomb lattice, where $i\in\{1,\dots,N_{x}\}$ and $j\in\{1,\dots,N_{y}\}$ label the unit cell and $s\in\{A,B\}$ labels the two sublattices, and $\sigma_{\mathbf{J}}^{\alpha}$ denotes the $\alpha=x,y,z$ Pauli matrix at site $\mathbf{J}$. The bonds of the lattice are partitioned into three sets – $x$, $y$, and $z$ bonds –according to their geometric orientations (see Fig. 2). We denote the corresponding Kitaev anisotropic exchange couplings by $\mathcal{J}_{x},\mathcal{J}_{y},$ and $\mathcal{J}_{z}$, respectively. The $\kappa$ term is a three-spin interaction acting on three neighboring sites denoted by $\langle\langle\mathbf{J},\mathbf{K},\mathbf{L}\rangle\rangle$, such that both $\mathbf{J}$ and $\mathbf{L}$ are nearest neighbors of site $\mathbf{K}$. The $\kappa$ term breaks time reversal symmetry, and opens a gap in the bulk that drives the system into the chiral non-Abelian phase Kitaev (2006).
Figure 2: The generalized KSL model with two system spins (black) and two bath spins (white) per unit cell. The bonds are colored according to orientation: x-blue, y-green, z-red. The five-spin term $\kappa$ drives the system into the chiral phase. A product of the unitaries $\mathcal{U}_{\mathbf{I},\mathbf{J}}$ along a path removes a pair of fermionic excitations from the bath. Instead of applying this non-local operation to the state, any subsequent operator acting on the system is conjugated by it.
@@ II.3 Preparation Protocol
    The $\sigma$ spins are brought to a flux-free state, $W_{i,j}=1$. This can be done by measuring all the flux operators (which are mutually commuting), and annihilating the detected fluxes in pairs by applying appropriate unitary operators Kalinowski et al. (2023) (see Appendix A for details).
@@
    The system is evolved from time $t_{n}$ to $t_{n+1}=t_{n}+T$ with the Hamiltonian $H(t-t_{n})$, where $H(t)$ is given in Eq. (2). Specifically, we set the effective Zeeman fields to $B_{\mathbf{J}}(t)=B(t-t_{n})\tau_{\mathbf{J}}(t_{n})$ and the “system-bath couplings” to $g_{\mathbf{J}}(t)=g(t-t_{n})$, with the functions $g(t)$, $B(t)$ as shown in Fig. 3 and given explicitly in Appendix B.
@@
Figure 3: Adiabatic evolution of the time dependent couplings $g$ and $B$ of the control Hamiltonian [Eq. (4)]. Explicit expressions are given in Appendix B.
Importantly, the flux operators $W_{i,j}$ commute with $H(t)$. Ideally, they therefore retain their initial values, $W_{i,j}=1$. In the presence of decoherence on a real quantum simulator, $W_{i,j}$ may flip during the cycle. This may require returning to step 1 between cycles, where the flux operators are measured again and unitary operators are applied to correct flipped fluxes, as discussed in Sec. III.2 and Appendix A.
@@ II.4 Performance analysis
$\dot{\varepsilon}(t)=-C\left[\varepsilon(t)-\varepsilon_{s}\right],$ (11)
@@
where $C>0$ is the cooling rate, and $\varepsilon_{s}$ is the steady state energy density. The quantities $C$ and $\varepsilon_{s}$ depend on the protocol parameters, such as $T$, $B_{0}$ and $g_{1}$, and in a realistic noisy quantum simulator, also on the noise rate; in the perfectly adiabatic, noiseless case, $\varepsilon_{s}\rightarrow\varepsilon_{0}$, the ground state energy density. Solving Eq. (11), we obtain $\varepsilon(t)=\varepsilon_{s}+[\varepsilon(0)-\varepsilon_{s}]e^{-Ct}$. In contrast, in a simple adiabatic cooling protocol where only pairs of fermionic excitations can be transferred to the bath, $\dot{\varepsilon}\propto-(\varepsilon-\varepsilon_{s})^{2}$, resulting in a power-law approach to the steady state, $\varepsilon-\varepsilon_{s}\sim 1/t$.
@@
The fact that the fermionic Hamiltonian (8) is quadratic allows us to analyze the cooling dynamics in a basis in which the excitations are decoupled. Moreover, within the flux-free sector, and assuming that the system is prepared initially such that $\tau^{z}_{\mathbf{J}}=1$ and hence the signs of the $B_{\mathbf{J}}$’s are spatially uniform, the Hamiltonian can be brought into a translationally invariant form by a unitary transformation. This allows an explicit analysis of the performance of the protocol in momentum space, which we present in Appendix C. The analysis verifies the exponential convergence to the steady state with the number of cycles performed, as anticipated above. Moreover, we find that in the ideal (noiseless) case, the steady state energy $\varepsilon_{s}$ depends only on the adiabaticity of the protocol with respect to the system’s gap, $E_{\text{gap}}$, whereas the cooling rate $C$ depends also on the the adiabaticity with respect to the avoided crossing gaps encountered during the time evolution with $H(t)$. In particular, in the limit where $T$ is large compared to $1/E_{\text{gap}}$, the ground state can be approached with an exponential accuracy.
@@
While these results are demonstrated for the solvable model $\widetilde{H}_{\text{KSL}}$, which can be mapped to free fermions, we expect them to hold more generally. For example, the performance is not expected to be parametrically affected if quartic interactions between the system fermions are added to Eq. (3). The application of our protocol to general interacting fermionic Hamiltonians is discussed further in Sec. IV.
III Numerical simulations
III.1 Performance without noise
@@
We simulate the fermion cooling algorithm numerically with and without noise using the efficient procedure described in Appendix D which relies on the model being one of free fermions. In all simulations we choose $\mathcal{J}_{x}=\mathcal{J}_{y}=\mathcal{J}_{z}=\mathcal{J}=1$, $\kappa=1$, in the chiral phase, and use protocol parameters $B_{0}=7$, $g_{1}=0.5$. Periodic boundary conditions are used unless otherwise stated.
@@
Fig. 4 shows $\nu^{\rm{sys}}$ (solid lines) and $\nu^{\rm{bath}}$ (dashed lines) at the end of a single cooling cycle before performing the measurement of the $\tau$ spins, as a function of the sweep time, $T$. As the sweep duration is increased, the Chern number of the system approaches $1$ and that of the bath approaches $-1$, even after a single cycle. The fact that our protocol succeeds in preparing a fermionic chiral ground state starting from a topologically trivial state by a finite depth local unitary circuit is due to the chiral state possessing “invertible” topological order; i.e., stacking two systems with opposite Chern numbers yields a trivial state. Thus, the topological state of the system’s fermions and its inverse (in the bath) can be prepared starting from a trivial state  [^1] Note, however, that overall, the spin system is in a non-invertible topologically ordered state, since the chiral phase of the KSL has fractional (non-Abelian) statistics..
Figure 4: A single cooling cycle is performed on a KSL system of size $85\times 85$ with $\mathcal{J}=1,\kappa=1$ starting from a randomly initialized state in the flux-free sector, using protocol parameters $g_{1}=0.5$, $B_{0}=7$. (a) A finite size version of the spectral Chern number (12) of the system (solid lines) and the bath (dashed lines) before performing the reset is shown as a function of the cycle duration $T$. The inset shows the energy density vs. $T$. In the adiabatic limit, the system approaches the ground state, the Chern number of the system approaches $1$ and that of the bath approaches $-1$ in a single cycle as illustrated in (b).
@@
III.2 Performance with noise
In this section, we numerically test the performance of the proposed cooling algorithm in the presence of noise. For simplicity, the noise is modeled as a uniform depolarizing channel acting on all $\sigma$ and $\tau$ qubits at some rate $\zeta$ of errors per cooling cycle per qubit. This noise is simulated stochastically as described in Appendix D.
Importantly, the fluxes of the gauge field may be excited by the action of noise on the system. In order to overcome this, one should periodically measure and fix these fluxes, while performing our proposed algorithm for cooling the excitations of the fermions $\{c_{\bf J}^{z}\}$. Fixing the observed fluxes can be done in an analogous way to the action of error correction in the surface code Dennis et al. (2002) by annihilating them in pairs, as explained in Appendix A. When flux excitations are removed (by applying local spin operations), fermionic excitations may be created, and these are then cooled by subsequent cycles of the cooling protocol. Using the protocol, the energy density of the steady state should be proportional to the noise rate, although there is a finite density of flux excitations in the steady state.
@@ III.2 Performance with noise
Figure 5: A KSL system of size $6\times 6$ unit cells initialized in its ground state is cooled in the presence of a decoherence rate of $\zeta=10^{-3}$ errors per cycle per qubit. The cycle duration is set to $T=20$ and $N_{t}=800$ Trotter steps are used per cycle. The energy density is plotted at the end of each cycle. The case where fluxes are corrected every cycle is shown in blue and the case where fluxes are left uncorrected is in orange.
@@
Figure 6: Energy density of a KSL model of size $10\times 10$ unit cells in the steady state reached after many cooling cycles as a function of the rate of depolarizing noise. The energies are averaged over $2000$ cycles per data point. The cycle duration is set to $T=10$ and $N_{t}=400$ Trotter steps are used per cycle. Post selection is applied with respect to the weight of decoded Pauli correction applied to fix the detected flux excitations, exponentially averaged over recent cycles $\bar{w}_{n}$. Different colors correspond to different fractions of cycles accepted as ranked by this criterion.
IV Extension to arbitrary fermionic Hamiltonians
Our protocol for preparing the ground state of the KSL model can be generalized as a method for preparing the ground state of an arbitrary (interacting) fermionic Hamiltonian in 2D. Given some local target Hamiltonian acting on the system Majorana fermions $c_{\mathbf{J}}^{z}$, we introduce the gauge fields $u_{\mathbf{I},\mathbf{J}}$ artificially and modify the target Hamiltonian such that the $c_{\mathbf{J}}^{z}$ fermions become charged under the gauge field. Namely, we modify each hopping term between site $\mathbf{I}$ and site $\mathbf{J}$ in the target Hamiltonian by the product of the gauge fields on a path connecting the two sites. We then introduce the bath fermions $c_{\mathbf{J}}^{x},c_{\mathbf{J}}^{y}$, and use Eq. (6) to map the fermionic Hamiltonian (including the hopping between the system and bath fermions, used for cooling the system) onto a bosonic local Hamiltonian acting on the $\sigma$ and $\tau$ spins. The ground state of the Hamiltonian is prepared according to the protocol above and this state is related to the ground state of the target Hamiltonian by the known gauge transformation as discussed in Sec. II.

### v2

@@ Abstract
Simulating many-body systems is one of the most promising applications of near-term quantum computers. An important open question is how to efficiently prepare the ground states of arbitrary fermionic Hamiltonians, especially those with nontrivial topology. Here, we propose an efficient protocol for preparing low-energy states of fermionic Hamiltonians on a noisy bosonic quantum simulator by adiabatic cooling using a simulated bath. We arrange the couplings such that the simulated system and bath together obtain a local fermionic description in which fermionic excitations can be extracted individually, via coherent hopping to the bath, rather than in pairs as would otherwise be required by fermion parity conservation. This approach thus achieves a linear (rather than quadratic) scaling of the cooling rate vs. excitation density at low densities. We show that certain topological phases such as the chiral (non-Abelian) phase of the Kitaev honeycomb model can be prepared efficiently using our protocol. Our protocol performs favorably in the presence of noise, making it suitable for execution on near-term quantum devices.
@@ I Introduction
Simulating quantum many-body systems, both in and out of equilibrium, is one of most promising applications of near-term quantum computers Cirac and Zoller (2012); Georgescu et al. (2014). In such tasks, quantum computers offer a vast advantage over their classical counterparts, by virtue of their ability to store the many-body wave functions and perform unitary gates that implement physical time evolution according to arbitrary Hamiltonians, at a polynomial cost in the system size. Such simulations may have practical utility in many fields, such as materials science, high-energy physics, and quantum chemistry. Preparation of a low-energy state of a many-body Hamiltonian is an important subroutine for these simulations. Doing this efficiently for arbitrary Hamiltonians is a highly non-trivial task, and is one of the major challenges in the field.
@@
The preparation scheme consists of two steps. In the first step, the system is prepared in the flux-free sector by measuring the flux operators through all plaquettes, and applying unitary operators that correct the fluxes that are detected Dennis et al. (2002). The second step is to prepare the fermionic ground state, whose Chern number is non-zero. The key idea is to introduce auxiliary “bath” degrees of freedom that simulate an effective fermionic bath, coupled to the system’s emergent fermions. Both species of fermions are coupled to the same $\mathbb{Z}_{2}$ gauge field, allowing single-fermion hopping between them. We provide an explicit encoding of the system and bath fermions into the physical bosonic qubits, generalizing the one-dimensional construction of Ref. Kishony et al. (2023) to two dimensions. (By bosonic qubits, we mean that operators belonging to different qubits commute with each other. Qubits can be thought of as hard-core bosons.) The bath fermions are then used to extract energy and entropy from the system, by performing simulated cyclic adiabatic cooling Matthies et al. (2022); Kishony et al. (2023) followed by measurements of the auxiliary qubits. The steps of the protocol are illustrated in Fig. 1.
Figure 1: Scheme for simulated cooling of a fermionic system using bosonic qubits. The system’s fermionic degrees of freedom (black) together with a fermionic “bath” (white) are encoded using bosonic qubits such that the two are charged under the same $\mathbb{Z}_{2}$ gauge field. (a) The system is brought to a flux-free state by measuring the flux operators and annihilating defects in pairs by applying a string of spin operators. (b) The system is evolved unitarily with the time-dependent Hamiltonian in Eq. (2), allowing single fermionic excitations to hop coherently from the system to the bath. (c) At the end of the cooling cycle, the hopping between the system and bath sites is turned off. The location of the fermionic excitations in the bath is measured and the measurement outcomes are used to determine the Hamiltonian for the next cycle.
@@
II Results
II.1 Preparing a chiral spin liquid
II.1.1 Model
@@
Here, $\mathbf{J}=(i,j,s)$ denotes the sites of the honeycomb lattice, where $i\in\{1,\dots,N_{x}\}$ and $j\in\{1,\dots,N_{y}\}$ label the unit cell and $s\in\{A,B\}$ labels the two sublattices, and $\sigma_{\mathbf{J}}^{\alpha}$ denotes the $\alpha=x,y,z$ Pauli matrix at site $\mathbf{J}$. The bonds of the lattice are partitioned into three sets – $x$, $y$, and $z$ bonds –according to their geometric orientations (see Fig. 2). We denote the corresponding Kitaev anisotropic exchange couplings by $\mathcal{J}_{x},\mathcal{J}_{y},$ and $\mathcal{J}_{z}$, respectively. The $\kappa$ term is a three-spin interaction acting on three neighboring sites denoted by $\langle\langle\mathbf{J},\mathbf{K},\mathbf{L}\rangle\rangle$, such that both $\mathbf{J}$ and $\mathbf{K}$ are nearest neighbors of site $\mathbf{L}$. The $\kappa$ term breaks time reversal symmetry, and opens a gap in the bulk that drives the system into the chiral non-Abelian phase Kitaev (2006).
@@
Figure 2: The generalized KSL model. Each unit cell consists of two system spins (black) and two bath spins (white). The bonds are colored according to orientation: x-blue, y-green, z-red. The five-spin term $\kappa$ drives the system into the chiral phase. A product of the unitaries $\mathcal{U}_{\mathbf{I},\mathbf{J}}$ along a path removes a pair of fermionic excitations from the bath. Instead of applying this non-local operation to the state, any subsequent operator acting on the system is conjugated by it.
@@ II.1.3 Preparation Protocol
    The $\sigma$ spins are brought to a flux-free state, $W_{i,j}=1$. This can be done by measuring all the flux operators (which are mutually commuting), and annihilating the detected fluxes in pairs by applying appropriate unitary operators Kalinowski et al. (2023) (see Methods subsection “Removing fluxes in 2D” for details).
@@
    The system is evolved from time $t_{n}$ to $t_{n+1}=t_{n}+T$ with the Hamiltonian $H(t-t_{n})$, where $H(t)$ is given in Eq. (2). Specifically, we set the effective Zeeman fields to $B_{\mathbf{J}}(t)=B(t-t_{n})\tau_{\mathbf{J}}(t_{n})$ and the “system-bath couplings” to $g_{\mathbf{J}}(t)=g(t-t_{n})$, with the functions $g(t)$, $B(t)$ as shown in Fig. 3 and given explicitly in Methods subsection “Smooth evolution of time dependent couplings”.
@@
Figure 3: Adiabatic evolution of the time dependent couplings $g$ and $B$. The time dependence of the couplings $g$ and $B$ of the control Hamiltonian [Eq. (4)]. Explicit expressions are given in Methods subsection “Smooth evolution of time dependent couplings”.
Importantly, the flux operators $W_{i,j}$ commute with $H(t)$. Ideally, they therefore retain their initial values, $W_{i,j}=1$. In the presence of decoherence on a real quantum simulator, $W_{i,j}$ may flip during the cycle. This may require returning to step 1 between cycles, where the flux operators are measured again and unitary operators are applied to correct flipped fluxes, as discussed in Sec. “Performance with noise” and in Methods subsection “Removing fluxes in 2D”.
@@ II.1.4 Performance analysis
$\dot{\varepsilon}(t)=-C\left[\varepsilon(t)-\varepsilon_{\rm s}\right],$ (11)
@@
where $C>0$ is the cooling rate, and $\varepsilon_{\rm s}$ is the steady state energy density. The quantities $C$ and $\varepsilon_{\rm s}$ depend on the protocol parameters, such as $T$, $B_{0}$ and $g_{1}$, and in a realistic noisy quantum simulator, also on the noise rate; in the perfectly adiabatic, noiseless case, $\varepsilon_{\rm s}\rightarrow\varepsilon_{0}$, the ground state energy density. Solving Eq. (11), we obtain $\varepsilon(t)=\varepsilon_{\rm s}+[\varepsilon(0)-\varepsilon_{\rm s}]e^{-Ct}$. In contrast, in a simple adiabatic cooling protocol where only pairs of fermionic excitations can be transferred to the bath, $\dot{\varepsilon}\propto-(\varepsilon-\varepsilon_{\rm s})^{2}$, resulting in a power-law approach to the steady state, $\varepsilon-\varepsilon_{\rm s}\sim 1/t$.
@@
The fact that the fermionic Hamiltonian (8) is quadratic allows us to analyze the cooling dynamics in a basis in which the excitations are decoupled. Moreover, within the flux-free sector, and assuming that the system is prepared initially such that $\tau^{z}_{\mathbf{J}}=1$ and hence the signs of the $B_{\mathbf{J}}$’s are spatially uniform, the Hamiltonian can be brought into a translationally invariant form by a unitary transformation. This allows an explicit analysis of the performance of the protocol in momentum space, which we present in Supplementary Note 1, Supplementary Fig. 1 and Supplementary Fig. 2. The analysis verifies the exponential convergence to the steady state with the number of cycles performed, as anticipated above. Moreover, we find that in the ideal (noiseless) case, the steady state energy $\varepsilon_{\rm s}$ depends only on the adiabaticity of the protocol with respect to the system’s gap, $E_{\text{gap}}$, whereas the cooling rate $C$ depends also on the the adiabaticity with respect to the avoided crossing gaps encountered during the time evolution with $H(t)$. In particular, in the limit where $T$ is large compared to $1/E_{\text{gap}}$, the ground state can be approached with an exponential accuracy.
@@
While these results are demonstrated for the solvable model $\widetilde{H}_{\text{KSL}}$, which can be mapped to free fermions, we expect them to hold more generally. For example, the performance is not expected to be parametrically affected if quartic interactions between the system fermions are added to Eq. (3). In the presence of interactions, excitations retain a finite overlap with single fermions since the non-interacting model is gapped. The application of our protocol to general interacting fermionic Hamiltonians is discussed further in Sec. “Extension to arbitrary fermionic Hamiltonians”.
II.2 Numerical simulations
II.2.1 Performance without noise
@@
We simulate the fermion cooling algorithm numerically with and without noise using the efficient procedure described in Methods subsection “Single-particle density matrix simulation of free fermions” which relies on the model being one of free fermions. In all simulations we choose $\mathcal{J}_{x}=\mathcal{J}_{y}=\mathcal{J}_{z}=\mathcal{J}=1$, $\kappa=1$, in the chiral phase, and use protocol parameters $B_{0}=7$, $g_{1}=0.5$. Periodic boundary conditions are used unless otherwise stated.
@@
Fig. 4 shows $\nu^{\rm{sys}}$ (solid lines) and $\nu^{\rm{bath}}$ (dashed lines) at the end of a single cooling cycle before performing the measurement of the $\tau$ spins, as a function of the sweep time, $T$. As the sweep duration is increased, the Chern number of the system approaches $1$ and that of the bath approaches $-1$, even after a single cycle. The fact that our protocol succeeds in preparing a fermionic chiral ground state starting from a topologically trivial state by a finite depth local unitary circuit is due to the chiral state possessing “invertible” topological order; i.e., stacking two systems with opposite Chern numbers yields a trivial state. Thus, the topological state of the system’s fermions and its inverse (in the bath) can be prepared starting from a trivial state. Note, however, that overall, the spin system is in a non-invertible topologically ordered state, since the chiral phase of the KSL has fractional (non-Abelian) statistics.
Figure 4: A single cooling cycle performed on the KSL system. A KSL system of size $85\times 85$ with $\mathcal{J}=1,\kappa=1$ starting from a randomly initialized state in the flux-free sector is cooled for a single cycle using protocol parameters $g_{1}=0.5$, $B_{0}=7$. (a) A finite size version of the spectral Chern number (12) of the system (solid lines) and the bath (dashed lines) before performing the reset is shown as a function of the cycle duration $T$. The inset shows the energy density vs. $T$. In the adiabatic limit, the system approaches the ground state, the Chern number of the system approaches $1$ and that of the bath approaches $-1$ in a single cycle as illustrated in (b).
@@
II.2.2 Performance with noise
In this section, we numerically test the performance of the proposed cooling algorithm in the presence of noise. For simplicity, the noise is modeled as a uniform depolarizing channel acting on all $\sigma$ and $\tau$ qubits at some rate $\zeta$ of errors per cooling cycle per qubit. This noise is simulated stochastically as described in Methods subsection “Single-particle density matrix simulation of free fermions”.
Importantly, the fluxes of the gauge field may be excited by the action of noise on the system. In order to overcome this, one should periodically measure and fix these fluxes, while performing our proposed algorithm for cooling the excitations of the fermions $\{c_{\bf J}^{z}\}$. Fixing the observed fluxes can be done in an analogous way to the action of error correction in the surface code Dennis et al. (2002) by annihilating them in pairs, as explained in Methods subsection “Removing fluxes in 2D”. When flux excitations are removed (by applying local spin operations), fermionic excitations may be created, and these are then cooled by subsequent cycles of the cooling protocol. Using the protocol, the energy density of the steady state should be proportional to the noise rate, although there is a finite density of flux excitations in the steady state.
@@ II.2.2 Performance with noise
Figure 5: Cooling in the presence of noise. A KSL system of size $6\times 6$ unit cells initialized in its ground state is cooled in the presence of a decoherence rate of $\zeta=10^{-3}$ errors per cycle per qubit. The cycle duration is set to $T=20$ and $N_{t}=800$ Trotter steps are used per cycle. The energy density is plotted at the end of each cycle. The case where fluxes are corrected every cycle is shown in blue and the case where fluxes are left uncorrected is in orange.
@@
Figure 6: Steady state in the presence of noise. The energy density of a KSL model of size $10\times 10$ unit cells in the steady state reached after many cooling cycles as a function of the rate of depolarizing noise. The energies are averaged over $2000$ cycles per data point and error bars represent the standard error of the mean determined for an effective sample size of $2000/(2n_{0})$ taking into account exponentially decaying correlations over approximately $n_{0}=3$ cycles. The cycle duration is set to $T=10$ and $N_{t}=400$ Trotter steps are used per cycle. The flux degrees of freedom are monitored and corrected at the end of each cycle. Post selection is performed with respect to the weight of decoded Pauli correction applied to fix the detected flux excitations, exponentially averaged over recent cycles $\bar{w}_{n}$. Cycles requiring Pauli corrections of high weight are omitted. Different colors correspond to different fractions of cycles accepted as ranked by this criterion. The data is fit by linear functions (dashed lines).
@@
II.3 Extension to arbitrary fermionic Hamiltonians
Our protocol for preparing the ground state of the KSL model can be generalized as a method for preparing the ground state of an arbitrary (interacting) fermionic Hamiltonian in 2D. Given some local target Hamiltonian acting on the system Majorana fermions $c_{\mathbf{J}}^{z}$, we introduce the gauge fields $u_{\mathbf{I},\mathbf{J}}$ artificially and modify the target Hamiltonian such that the $c_{\mathbf{J}}^{z}$ fermions become charged under the gauge field. Namely, we modify each hopping term between site $\mathbf{I}$ and site $\mathbf{J}$ in the target Hamiltonian by the product of the gauge fields on a path connecting the two sites. We then introduce the bath fermions $c_{\mathbf{J}}^{x},c_{\mathbf{J}}^{y}$, and use Eq. (6) to map the fermionic Hamiltonian (including the hopping between the system and bath fermions, used for cooling the system) onto a bosonic local Hamiltonian acting on the $\sigma$ and $\tau$ spins. The ground state of the Hamiltonian is prepared according to the protocol above and this state is related to the ground state of the target Hamiltonian by the known gauge transformation as discussed in Sec. “Preparing a chiral spin liquid”.